At the beginning of 2010 Monique's car was worth $10,000, but the value of her car decreases exponentially. She notices that the value of her car decreases by 20% every 3 years.
a. What is the 3-year growth factor for the value of Monique's car?
b. What is the 1-year growth factor for the value of Monique's car?
c. Write a function f that determines the value of Monique's car (in dollars) in terms of the number of years t since the beginning of 2010.

Answer by Theo(13342)

f = p * (1 + r) ^ n

f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

that's the general formula for the future value of something when compound interest is concerned.

compound interest is a form of exponential growth or decay.

if the car cost 10,000 and decreases 20% in value every 3 years, then:

you can solve as follows:

time period is every 3 years.
interest rate is interest rate per every 3 years.

f = p * (1 + r) ^ n becomes f = 10,000 * (1 - .20) ^ n

simplify this to get f = 10,000 * .80 ^ n

after 3 years, the value of the car is 10,000 * .8 ^ 1 = 8000.
after 6 years, the value of the car is 10,000 * .8 ^ 2 = 6400.
after 9 years, the value of the car is 10,000 * .8 ^ 3 = 5121.536

the average annual growth rate would be calculated as follows:

find the future value after one 3 year period.

assuming the present value is one, the formula would become.

f = 1 * (1 - .2) ^ 1 = 1 * .8 = .8

so, the present value becomes .8 after 3 years.

to find the annual growth rate, you would do the following.

f = p * (1 + r) ^ n becomes .8 = 1 * (1 + r) ^ 3

divide both sides of the equation by 1 to get .8 = (1 + r) ^ 3

take the cube root of both sides of the eqution to get .8 ^ (1/3) = 1 + r

subtract 1 from both sides of the equation to get .8 ^ (1/3) - 1 = r

solve for r to get r = -.0716822333.

that's the annual growth rate.

to confirm, use this rate in the original equation of .8 = 1 * (1 + r) ^ 3

you will get .8 = 1 * (1 - .0716822333) ^ 3

simplify to get .8 = .8

if you graph this equation, it will look like this.

$$$

from the graph, you can see that the loss in value is 20% every 3 years.

in year 0, the value is 10,000
in year 3, the value is 8000
in year 6, the value is 6400
in year 9, the value is 5120

the every 3 year growth rate is (1 - .2) = .8

the every year growth rate is (1 - .0716822333) = .9283177667